A metal with body centered cubic (bcc) structure shows the first (i.e. smallest angle) diffraction peak at Bragg angle of θ=30∘. The wavelength of X-ray used is 2.1Å. The volume of primitive unit cell of the metal is :


26.2(Å3)




13.1(Å3)




9.3(Å3)




4.6(Å3)


1
Expert's answer
2021-10-19T15:15:40-0400

Determine the length of the side of the primitive unit cell with Bragg's law:


"\lambda=2d\sin\theta,\\\space\\
d=\frac\lambda{2\sin\theta},\\\space\\
V=d^3=\bigg(\frac\lambda{2\sin\theta}\bigg)^3=9.3\text{ Å}^3."